What is the electric potential at the center of a square of side length 2.0 m with alternating + 3C, –2C, +3C, and –2C charges placed at its vertices


1
Expert's answer
2021-03-23T11:15:03-0400

Answer

net charge

"q_{net}=3+3-2-2=2C"


Distance of all the charge from center

="\frac{\sqrt{2^2+2^2}}{2}=1.4m"


So potential at center

"V=\frac{kq}{r^2}=\frac{(9\times10^9) (2) }{1.4^2}=9.2\times10^9V"